OpenStudy (anonymous):

Find the derivative of f(x) = 6/x at x = -2. help

OpenStudy (ipwnbunnies):

Rewrite f(x) so x is in the numerator. It'll be easier to take the derivative.

OpenStudy (anonymous):

how do i do that

zepdrix (zepdrix):

\[\Large\bf\sf f(x)\quad=\quad \frac{6}{x}\]Use this rule of exponents, as bunny mentioned. \[\Large\bf\sf \frac{1}{x}\quad=\quad x^{-1}\]

OpenStudy (anonymous):

so x^-6

zepdrix (zepdrix):

Nooo you silly billy!\[\Large\bf\sf \frac{6}{x}\quad=\quad 6\frac{1}{x}\quad=\quad 6x^{-1}\]

zepdrix (zepdrix):

Do you understand how I was able to pull the 6 out of the numerator like that?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

but i don't get what to do after that

zepdrix (zepdrix):

Power ruuuuule \c:/

zepdrix (zepdrix):

Remember your power rule?

OpenStudy (ipwnbunnies):

POWUH RULE IS THE BESTTT RULE.

zepdrix (zepdrix):

One Power to Rule Them All!! \c:/

OpenStudy (anonymous):

my guess is you are supposed to do this by hand, not power rule

OpenStudy (anonymous):

do you put it under a radical

OpenStudy (anonymous):

my guess is you have to compute \[\lim_{x\to -2}\frac{f(x)-f(-2)}{x-(-2)}\]

OpenStudy (anonymous):

does that look familiar?

zepdrix (zepdrix):

How you not learned the shortcut rules at this point? Still using the limit definition of derivative? A little clarity would help ^^

OpenStudy (anonymous):

would the answer be 3/2?

OpenStudy (anonymous):

the derivative of \(\frac{1}{x}\) is \(-\frac{1}{x^2}\) so the derivative of \(\frac{6}{x}\) is \(-\frac{6}{x^2}\)

OpenStudy (anonymous):

so i was right?

OpenStudy (anonymous):

if \(x=-2\) you get \(-\frac{6}{(-2)^2}\) or \(-\frac{3}{2}\)

OpenStudy (anonymous):

you were off by a minus sign

OpenStudy (anonymous):

ohh.. darn. well thank you for helping :)